Mir Status Report 18 September 1997

NASA MIR Program, Mir 24 / NASA 5 Status
Date: Thursday, September 18, 1997 // MD 44/126

Mir Station System/Subsystem Activities

Gyrodynes #2, 5, and 6 on Kvant-2 were spun up last night, but #1 did not spin up, and #6 was spinning up more slowly than the others. This brings the number of active units to nine, although #6 is not yet in the control loop. Work continues on Kvant-2 #1 Gyrodyne, and Mr. Antoshechkin informed us that the crew would replace the main axis drive control unit for Kvant-1 Gyrodyne #3 and attempt to spin it up later today.

The Kvant-1 Elektron was reactivated at DMT 16:33 yesterday and is running fine.

Core module external loop KOX-1H was switched off at DMT 21:27 last night due to the dewpoint being reached on the Antares coldplate on the joint KOX-1B/VGK loop. Both Altair passes for today were cancelled due to the dewpoint concern. Mr. Antoshechkin told us that one fix would be to raise the setpoint of the KOX-1H loop to bring the Antares coldplate above the dewpoint, but this would also increase the temperature of the coolant flowing to Elektron, so they cannot exercise this option.

Mir Payload Activities
Greenhouse Daily Ops
Beetle Ventilation and Power Status Check

OTHER ACTIVITIES
Cognilab
Mir Cable Installation

Greenhouse - yesterday, Mike reported that the seeds collected were different from the first planting's. The seeds in Cuvette 2 were less numerous. They were fairly large in size. The seed color is light brown as opposed to a dark blackish brown color of the seeds collected previously. The seeds recently collected compared much more favorably with the Earth seeds which are still slightly bigger than the space seeds that were just planted. The Cuvette 1 seeds were more similar to the seeds planted in the first planting. The siliques in Cuvette 1 produced more seeds than the ones in Cuvette 2. The seeds from Cuvette 1 are somewhat flat, oval, and dark green in color.

Experiment Hardware

Priroda Module - It is currently anticipated that the Priroda module power up will occur tomorrow afternoon. Unfortunately it will occur during a period in which no ground telemetry will occur. Ground telemetry is required for the ground controllers responsible for the Priroda module to monitor the system status and declare the systems fully operational. It is hoped that some Priroda operations, such as PUP checkout, can be scheduled for Saturday.

SMP/Medical -- Mike will begin doing two treadmill sessions per day as per the nominal Russian end-of-mission plan.

Environmental Parameters (Previous 24 hour period):
Parameter Min. Max. Nominal Range
Pressure (mmHg) Base Block 617 620 660 - 850
ppO2 (mmHg) 151 154 140 - 200
ppCO2 (mmHg) 4.1 4.7 up to 8.0
ppH20 (mmHg) 13.0 14.1 8 - 14
Humidity (percentage) 55 59 30 - 75
Temperature (xC) base block (BB) 24.7 25.7 18 - 28

Communications Summary / A/G with Consultants Group
09:44-09:53 CGEL; EVA video; prepacking
11:20-11:25 Prepacking
16:16-16:21 MIPS; Priroda lack of space; prepacking
19:01-19:06 PAO Speech

  • submit to reddit


Loading



MILITARY SPACE USA, 11TH JUNE TO 12TH JUNE 2019, LOS ANGELES, USA
49th International Conference on Environmental Systems - ICES 2019




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on September 18, 1997 8:45 PM.

Mir Status Report 17 September 1997 was the previous entry in this blog.

Mir Status Report 19 September 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.